OpenStudy (anonymous):

Solve using substitution method x + 6y = 1 2x + 5y = 9

OpenStudy (anonymous):

Top: x= 1-6y Bottom: 2(1-6y) +5y=9 2-12y + 5y=9 2-17y=9 -17y=7 y=7/-17

OpenStudy (anonymous):

I got y=-1

OpenStudy (anonymous):

and x=7?

OpenStudy (jim766):

yeah above it should say 2 - 7y = 9 not 2 - 17 = 9

OpenStudy (jim766):

right.....

OpenStudy (anonymous):

okay so the solution would be (7,-1) ?

OpenStudy (jim766):

right!

OpenStudy (anonymous):

Thank you(: